Please help me solve the equation
b^4+13b^2+36=0
Find all real and imaginery solutions to this problem

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

let a=b^2
a^2+13a+36=0
find factors of 36 which add up to 13
1 36
2 18
3 13
4 9
6 6
(a+9)*(a+4)=0
(b^2+9)*(b^2+4)=0
b^2=-9
b^2=-4
b=2i,-2i,3i,-3i

Answer by stanbon(75887)

b^4+13b^2+36=0
---
b^4 + 9b^2 + 4b^2 + 36 = 0
--------
b^2(b^2+9) + 4(b^2+9) = 0
--------
(b^2+9)(b^2+4) = 0
----
(b+3i)(b-3i)(b+2i)(b-2i) = 0
----
b = 3i or b = -3i or b = 2i or b = -2i
================================================
Cheers,
Stan H.
================================================